# Nonmonotone subgradient methods based on a local descent lemma

### v1

@@ 4 A nonmonotone subgradient method for nonconvex minimization
Let us first briefly draw our attention to the optimality conditions for (P). If $\bar{x}\in\mathbb{R}^{n}$ is a local minimum, it follows by Fermat’s rule (see, e.g., [25, Proposition 1.30]) that
@@
We say that a point $\bar{x}$ satisfying (10) is stationary for problem (P).
@@
Remark 4.2 (On the choice of directions $d_{k}$).
@@
    $\lambda_{\max}(B_{k})\leq\lambda_{\texttt{M}},\quad\text{for all }k\in\mathbb{N}.$
  • Assumption (12), on the other hand, corresponds to imposing that for any bounded subsequence $(x_{k_{j}})_{j\in\mathbb{N}}$ generated by Algorithm 1, there exists a positive constant $\lambda_{\texttt{m}}$ such that
    $\lambda_{\texttt{m}}\leq\lambda_{\min}(B_{k_{j}}),\quad\text{for all }j\in\mathbb{N}.$
@@
    If $(x_{k_{j}})_{j\in\mathbb{N}}$ is a bounded subsequence of $(x_{k})_{k\in\mathbb{N}}$, then $\inf_{j\in\mathbb{N}}\tau_{k_{j}}>0$. In addition, we have that $\lim\limits_{j\to\infty}\|x_{k_{j}+1}-x_{k_{j}}\|=0$, $\lim\limits_{j\to\infty}\|w_{k_{j}}\|=0$, $\lim\limits_{j\to\infty}\|d_{k_{j}}\|=0$ and $\lim\limits_{j\to\infty}\varphi(x_{k_{j}})=\lim\limits_{k\to\infty}\varphi(x_{\ell(k)})$.
@@
We proceed by way of contradiction. Suppose there exists a subsequence $(\tau_{\nu_{i}})_{i\in\mathbb{N}}$ of $(\tau_{k_{j}})_{j\in\mathbb{N}}$ such that $\tau_{\nu_{i}}\to 0^{+}$ as $i\to\infty$. Given the boundedness of the sequence $(x_{k_{j}})_{j\in\mathbb{N}}$, we may, without loss of generality, assume that the corresponding subsequence $(x_{\nu_{i}})_{i\in\mathbb{N}}$ converges to some point $\bar{x}$. Moreover, since $\varphi$ is locally Lipschitz continuous and each $w_{\nu_{i}}$ belongs to $\partial\varphi(x_{\nu_{i}})$, it follows that the sequence $(w_{\nu_{i}})_{i\in\mathbb{N}}$ is bounded (see, e.g., [11, Proposition 2.1.2]). Hence, by possibly passing to a smaller subsequence without relabeling, we may assume that $w_{\nu_{i}}\to\bar{w}$ for some $\bar{w}\in\mathbb{R}^{n}$.
@@
which implies that the set of accumulation points of $(x_{k})_{k\in\mathbb{N}}$ is nonempty, closed, and connected (see, e.g., [16, Theorem 8.3.9]).
@@
Suppose that $x_{k_{j}}\to\bar{x}$ when $j\to\infty$. Using assertion (i), we have that $w_{k_{j}}\to 0$ as $j\to\infty$. Since $w_{k_{j}}\in\partial\varphi(x_{k_{j}})$ for all $j\in\mathbb{N}$, the closedness of the subdifferential mapping of locally Lipschitz functions (see, e.g., [11, Proposition 2.1.5]) implies that $0\in\partial\varphi(\bar{x})$.
Assertion (iv) follows from [16, Proposition 8.3.10], which ensures that the whole sequence $(x_{k})_{k\in\mathbb{N}}$ converges to a point $\bar{x}$. This point $\bar{x}$ is stationary due to assertion (iii). ∎
@@
Although Algorithm 1 and Theorem 4.5 are stated in terms of the Clarke subdifferential, we emphasize that the same conclusions of Theorem 4.5 remain valid when the sequence $(w_{k})_{k\in\mathbb{N}}$ is selected from the (Mordukhovich) limiting subdifferential of $\varphi$. This follows from the fact that the limiting subdifferential is contained within the Clarke subdifferential, ensuring that the descent condition (3) still holds. In this case, the convergence result guarantees stationarity in the sense of the Mordukhovich subdifferential, which is a stronger condition than the Clarke-type stationarity defined in (10).
@@ 6 Minimum sum-of-squares clustering problem
The essential steps for the implementation of SNSM for this problem are the selection of a Clarke subgradient of $\varphi_{MSC}$ in step 4 and the computation of an appropriate descent direction in step 5. The former has been discussed in [8, Section 4.2], where the authors provided a close expression for an element in the limiting subdifferential of the functions $\omega^{j}$. Accordingly, we propose in step 4 of the $k$-th iteration to choose an element of the form
@@
where $X_{k}:=(x_{k}^{1},\ldots,x_{k}^{\ell})$. Note that, in general, $w_{k}$ is not guaranteed to belong to the limiting subdifferential because of the possible failure of the subdifferential sum rule, but $w_{k}\in\partial\varphi_{MSC}(X_{k})$ (see [8, Proposition 4.1]). This choice allows for an efficient subgradient evaluation while leads to an adequate optimality condition.
@@
This selection allows us to define a differentiable function
${\varphi}^{k}_{MSC}(X):=\frac{1}{p}\sum_{j=1}^{p}\|x^{i_{k}(j)}-a^{j}\|^{2},$
whose value at the current iterate coincides with that of the original function; namely, ${\varphi}^{k}_{MSC}({X_{k}})={\varphi}_{MSC}({X_{k}})$. Now, note that the Hessian of ${\varphi}^{k}_{MSC}$ corresponds to a diagonal matrix with diagonal given block-wise by twice the number of times that each block variable appears in the sum, i.e.,
@@
Observe that $\nabla^{2}{\varphi}^{k}_{MSC}$ may be singular at some point if some of the indices are not active for any $j\in\{1,2,\ldots,p\}$. To overcome this fact, we add a regularization term $\alpha_{k}>0$, and select $d_{k}$ in step 5 of SNSM as
@@
$d_{k}=-\left(\left(\frac{p}{2q_{1}(k)+\alpha_{k}},\frac{p}{2q_{2}(k)+\alpha_{k}},\ldots,\frac{p}{2q_{\ell}(k)+\alpha_{k}}\right)\otimes\mathbbm{1}_{s}\right)\odot w_{k},$
@@ 7 Numerical experiments
In the following, we present some computational tests where we compare the performance of SNSM against other commonly used algorithms. Specifically, we test the Difference of Convex functions Algorithm (DCA) introduced in [34], the inertial Difference of Convex functions Algorithm (iDCA) presented in [15] (which includes an extrapolation term to the classical DCA), the Boosted Difference of Convex function Algorithm (BDCA) proposed in [5, 7] (which incorporates a linesearch to DCA), as well as the Regularized Coderivative-based damped Semi-Newton method (RCSN) from [6].
@@
For all the algorithms and experiments, we used the stopping criterion
@@
All the tests were ran on a computer with an ADM Ryxen 9 7950X 16-Core Processor, 4.50 GHz with 64GM RAM, under Windows 10 (64-bit).
@@ 7.1 Quadratic optimization with integer constraints
Table 1: For different values of the space dimension $n$ and different radii of the balls, we compared the output of SNSM against DCA, iDCA, BDCA and PNA for 100 random starting points. We counted the instances in which SNSM obtained a higher/lower objective value than the other algorithms
@@ 7.2 Minimum Sum-of-Squares Clustering Problem
Further properties of this formulation can be consulted in, e.g., [13, 27]. This allows us to compare the implementation of SNSM presented in Section 6 with that of DCA, iDCA, BDCA and RCSN over the previous decomposition, thereby extending the numerical study in [7, Section 5.1]
@@
On the other hand, observe that RCSN applied to (28) results again in a particular case of Algorithm 2 (with $m=0$) where the direction $d_{k}$ is taken as $d_{k}=-[\nabla^{2}g(X_{k})]^{-1}w_{k}=-w_{k}/2$. In contrast, we implemented SNSM as described in Section 6, with the search direction given by (30) and $\alpha_{k}:=10^{-3}$ for all $k\in\mathbb{N}$. Thus, we emphasize that RCSN does not coincide in this case with the monotone implementation of SNSM, unlike in our experiments in Section 7.1. For this reason, we included in our comparison both the monotone version of SNSM (with $m=0$) and its nonmonotone counterpart (with $m=5$).
@@
For each data set, we ran all algorithms from 10 different random starting points and reported averaged results on the number of iterations, number of function evaluations, running time and obtained function value in Table 3. The experimental results show that SNSM beat the other methods in every single aspect. Only RCSN obtained a lower average running time for the data set Birch2, but at the cost of reaching a significantly higher function value. Regarding the comparison between the monotone and nonmonotone versions of SNSM, none of them seems to be consistently better than the other for this particular problem. Nonetheless, the nonmonotone version generally required a smaller number of iterations and function evaluations (except for Letters), and always needed less running time.
Data sets Method # Iterations # f evaluations Time Function value
@@
ConfLongdemo DCA 1149 $2299$ 1065.5 0.25
iDCA 1126 $2252$ 1043.2 0.25
BDCA 403 $1740$ 414.8 0.23
RCSN 212 $724$ 218.8 0.23
SNSM $m=0$ 33 $78$ 61.4 0.19
@@
Letters DCA 395 $790$ 66.2 57.04
iDCA 356 $712$ 59.6 57.38
BDCA 170 $738$ 44.1 52.41
RCSN 113 $383$ 29.6 54.60

### v2

@@ 4 A nonmonotone subgradient method for nonconvex minimization
Let us first briefly draw our attention to the optimality conditions for (P). If $\bar{x}\in\mathbb{R}^{n}$ is a local minimum, it follows by Fermat’s rule (see, e.g., [28, Proposition 1.30]) that
@@
We say that a point $\bar{x}$ satisfying (10) is a stationary point for problem (P).
@@
Remark 4.2 (Ensuring Assumption 4.1).
@@
    $\lambda_{\texttt{m}}\leq\lambda_{\min}(B_{k}),\quad\text{for all }k\in\mathbb{N}.$
  • Assumption (12), on the other hand, is satisfied whenever for any bounded subsequence $(x_{k_{j}})_{j\in\mathbb{N}}$ generated by Algorithm 1, there exists a positive constant $\lambda_{\texttt{M}}$ such that
    $\lambda_{\max}(B_{k_{j}})\leq\lambda_{\texttt{M}},\quad\text{for all }j\in\mathbb{N}.$
Lastly, condition (13) can be easily guaranteed by keeping a constant memory parameter. In Algorithm 2 we provide a different self-adaptive possibility.
@@
    If $(x_{k_{j}})_{j\in\mathbb{N}}$ is a bounded subsequence of $(x_{k})_{k\in\mathbb{N}}$, then $\inf_{j\in\mathbb{N}}\tau_{k_{j}}>0$. Moreover, $\lim\limits_{j\to\infty}\|x_{k_{j}+1}-x_{k_{j}}\|\hskip-1.13809pt=\hskip-1.13809pt0$, $\lim\limits_{j\to\infty}\|w_{k_{j}}\|\hskip-1.13809pt=\hskip-1.13809pt0$, $\lim\limits_{j\to\infty}\|d_{k_{j}}\|\hskip-1.13809pt=\hskip-1.13809pt0$ and $\lim\limits_{j\to\infty}\varphi(x_{k_{j}})\hskip-1.13809pt=\hskip-1.13809pt\lim\limits_{k\to\infty}\varphi(x_{\ell(k)})$.
@@
We proceed by way of contradiction. Suppose there exists a subsequence $(\tau_{\nu_{i}})_{i\in\mathbb{N}}$ of $(\tau_{k_{j}})_{j\in\mathbb{N}}$ such that $\tau_{\nu_{i}}\to 0^{+}$ as $i\to\infty$. Given the boundedness of the sequence $(x_{k_{j}})_{j\in\mathbb{N}}$, we may, without loss of generality, assume that the corresponding subsequence $(x_{\nu_{i}})_{i\in\mathbb{N}}$ converges to some point $\bar{x}$. Moreover, since $\varphi$ is locally Lipschitz continuous and each $w_{\nu_{i}}$ belongs to $\partial\varphi(x_{\nu_{i}})$, it follows that the sequence $(w_{\nu_{i}})_{i\in\mathbb{N}}$ is bounded (see, e.g., [12, Proposition 2.1.2]). Hence, by possibly passing to a further subsequence without relabeling, we may assume that $w_{\nu_{i}}\to\bar{w}$ for some $\bar{w}\in\mathbb{R}^{n}$.
@@
which implies that the set of accumulation points of $(x_{k})_{k\in\mathbb{N}}$ is nonempty, closed, and connected (see, e.g., [19, Theorem 8.3.9]).
@@
Suppose that $x_{k_{j}}\to\bar{x}$ when $j\to\infty$. Using assertion (i), we have that $w_{k_{j}}\to 0$ as $j\to\infty$. Since $w_{k_{j}}\in\partial\varphi(x_{k_{j}})$ for all $j\in\mathbb{N}$, the closedness of the subdifferential mapping of locally Lipschitz functions (see, e.g., [12, Proposition 2.1.5]) implies that $0\in\partial\varphi(\bar{x})$.
Assertion (iv) follows from [19, Proposition 8.3.10], which ensures that the whole sequence $(x_{k})_{k\in\mathbb{N}}$ converges to a point $\bar{x}$. This point $\bar{x}$ is stationary due to assertion (iii). ∎
@@
(i) Although Algorithm 1 and Theorem 4.5 are stated in terms of the Clarke subdifferential, we emphasize that the same conclusions of Theorem 4.5 remain valid when the sequence $(w_{k})_{k\in\mathbb{N}}$ is selected from the (Mordukhovich) limiting subdifferential of $\varphi$. This follows from the fact that the limiting subdifferential is contained within the Clarke subdifferential, ensuring that the descent condition (3) still holds. In this case, the convergence result guarantees stationarity in the sense of the Mordukhovich subdifferential, which is a stronger condition than the Clarke-type stationarity defined in (10).
@@
(ii) While assumptions of the form of $\sup_{k\in\mathbb{N}}\tau_{k}<+\infty$, as required in Theorem 4.5, are common in the literature (see, e.g., [7, Theorem 3.11]), this condition can also be enforced by restricting the trial stepsize $\bar{\tau}_{k}$ to be bounded from above by a positive constant in step 5 of Algorithm 1. This approach was adopted, for instance, in [6, Algorithm 1]. Notably, imposing this upper bound was unnecessary in our numerical experiments, since the sequence remained bounded even under the self-adaptive rule introduced in the next section.
@@ 6 Minimum sum-of-squares clustering problem
The essential steps for the implementation of SNSM for this problem are the selection of a Clarke subgradient of $\varphi_{MSC}$ in step 4 and the computation of an appropriate descent direction in step 5. The former has been discussed in [7, Section 4.2], where the authors provided a close expression for an element in the limiting subdifferential of the functions $\omega^{j}$. Accordingly, we propose in step 4 of the $k$-th iteration to choose an element of the form
@@
where $X_{k}:=(x_{k}^{1},\ldots,x_{k}^{\ell})$. Note that, in general, $w_{k}$ is not guaranteed to belong to the limiting subdifferential because of the possible failure of the subdifferential sum rule, but $w_{k}\in\partial\varphi_{MSC}(X_{k})$ (see [7, Proposition 4.1]). This choice allows for an efficient subgradient evaluation while leads to an adequate optimality condition.
@@
For such indices, we define the function
${\varphi}^{k}_{MSC}(X):=\frac{1}{p}\sum_{j=1}^{p}\|x^{i_{k}(j)}-a^{j}\|^{2}.$
Since $i_{k}(j)$ is fixed for all $j\in\{1,\ldots,p\}$, this function is just a sum of squared norms and hence differentiable. Moreover, its value at the current iterate coincides with that of the original function; i.e., ${\varphi}^{k}_{MSC}({X_{k}})={\varphi}_{MSC}({X_{k}})$. Now, note that the Hessian of ${\varphi}^{k}_{MSC}$ corresponds to a diagonal matrix with diagonal given block-wise by twice the number of times that each block variable appears in the sum, i.e.,
@@
Observe that $\nabla^{2}{\varphi}^{k}_{MSC}$ may be singular at some point if some of the indices are not active for any $j\in\{1,2,\ldots,p\}$. To overcome this problem, we add a regularization term $\alpha_{k}>0$, and select $d_{k}$ in step 5 of SNSM as
@@
$d_{k}=-\left(\left(\frac{p}{2q_{1}(k)+p\alpha_{k}},\frac{p}{2q_{2}(k)+p\alpha_{k}},\ldots,\frac{p}{2q_{\ell}(k)+p\alpha_{k}}\right)\otimes\mathbbm{1}_{s}\right)\odot w_{k},$
@@ 7 Numerical experiments
In the following, we present some computational tests where we compare the performance of SNSM against other commonly used algorithms. Specifically, we test the Difference of Convex functions Algorithm (DCA) introduced in [40], the inertial Difference of Convex functions Algorithm (iDCA) presented in [17] (which includes an extrapolation term to the classical DCA), the Boosted Difference of Convex function Algorithm (BDCA) proposed in [5, 8] (which incorporates a line search to DCA), as well as the Regularized Coderivative-based damped Semi-Newton method (RCSN) from [6]. In addition, in Subsection 7.2.1, we evaluate the performance of SNSM on large data sets against large-scale clustering solvers, namely the K-means and Mini-Batch K-Means (MBKM) algorithms [39].
@@
For all the algorithms and experiments (with the exception of those in Subsection 7.2.1), we used the stopping criterion
@@
All the tests were run on a computer with an ADM Ryxen 9 7950X 16-Core Processor, 4.50 GHz with 64GM RAM, under Windows 10 (64-bit).
@@ 7.1 Quadratic optimization with integer constraints
Table 1: For different values of the space dimension $n$ and different radii of the balls, we compared the output of SNSM against DCA, iDCA, BDCA and PNA for 100 random starting points. Each entry of the table is given by $n_{1}/n_{2}$ where $n_{1}$ (respectively, $n_{2}$) is the number of intstances where SNSM obtained higher (respectively, lower) objective function value than the competing algorithm. Since ties may occur, $n_{1}$ and $n_{2}$ do not necessarily sum to the total number of 100 instances
@@
To further evaluate the scalability of the proposed method in higher dimensional settings, we conducted a second numerical experiment focusing on binary quadratic programming. We selected five large-scale instances [^1] Specifically, the files named bqp1000-5, bqp1000-6,…, bqp1000-9 from the library https://biqmac.aau.at/library/biq/beasley. (with $n=1000$) from the well-known OR-Library [10], which serve as a standard benchmark for NP-hard combinatorial optimization. For each instance, the algorithms were executed starting from 20 random initial vectors uniformly distributed in $0,1]^{n}$. For each initialization, we recorded the best objective function value achieved by each method across all tested radii. Figure 4 displays the performance profile of the objective values [18], which illustrates the percentage of problems for which a method’s objective value is within a factor $\varsigma$ of the best result found by any compared method. The numerical results demonstrate that SNSM outperforms the other methods. Specifically, SNSM achieved the best objective value in approximately 80% of the tested instances (as indicated by its value at $\varsigma=1$). Furthermore, its rapid ascent to the 1.0 level shows that it maintains a minimal deviation from the best-found value in the remaining instances, thereby exhibiting superior robustness.
Figure 4: Performance profiles for the objective function values on binary quadratic programming instances from the OR-Library
@@ 7.2 Minimum Sum-of-Squares Clustering Problem
Further properties of this formulation can be consulted in, e.g., [14, 30]. This allows us to compare the implementation of SNSM presented in Section 6 with that of DCA, iDCA, BDCA and RCSN over the previous decomposition, thereby extending the numerical study in [8, Section 5.1]
@@
On the other hand, observe that RCSN applied to (28) results again in a particular case of Algorithm 2 (with $m=0$) where the direction $d_{k}$ is taken as $d_{k}=-[\nabla^{2}g(X_{k})]^{-1}w_{k}=-w_{k}/2$. In contrast, we implemented SNSM as described in Section 6, with the search direction given by (30) and $\alpha_{k}:=10^{-3}$ for all $k\in\mathbb{N}$. We emphasize that RCSN does not coincide in this case with the monotone implementation of SNSM, unlike in our experiments in Section 7.1. Indeed, observe that the first component $g$ is just a sum of squared Euclidean norms and hence the Hessian is simply a multiple of the identity matrix, which leads to a simple scaling of the gradient descent direction and fails to capture the true curvature of the problem. For this reason, we included in our comparison both the monotone version of SNSM (with $m=0$) and its nonmonotone counterpart (with $m=5$).
@@
For each dataset, we ran all algorithms from 10 different random starting points and reported averaged results on the number of iterations, number of function evaluations, running time and obtained function value in Table 3. The experimental results show that SNSM beat the other methods in every single aspect. Only RCSN obtained a lower average running time for the dataset Birch2, but at the cost of reaching a significantly higher function value. Regarding the comparison between the monotone and nonmonotone versions of SNSM, none of them seems to be consistently better than the other for this particular problem.
Data set Method # Iterations # f evaluations Time Function value
@@
ConfLongdemo DCA 1149 $2299$ 41.8 0.25
iDCA 1126 $2252$ 40.8 0.25
BDCA 403 $1740$ 28.6 0.23
RCSN 218 $747$ 15.6 0.23
SNSM $m=0$ 43 $103$ 2.6 0.19
@@
Letters DCA 395 $790$ 3.7 57.04
iDCA 356 $712$ 3.2 57.38
BDCA 170 $738$ 3.0 52.41
RCSN 113 $383$ 2.0 54.60
@@
7.2.1 Comparison with large-scale clustering solvers
While SNSM was not primarily developed to compete with algorithms tailored for large datasets, we performed further numerical experiments to assess its scalability. Specifically, we compared our proposal against the standard K-means and Mini-Batch K-Means (MBKM) algorithms, as implemented in the scikit-learn Python library [32]. Because scikit-learn’s K-means and MBKM rely on highly optimized Cython and C-extensions, we implemented our algorithm using Numba’s Just-In-Time (JIT) compiler to guarantee a more fair runtime comparison. We also replaced our stopping criterion (31) by
$\frac{\|x_{k}-x_{k-1}\|}{\max\{\|x_{k-1}\|,1\}}\leq 10^{-4},$ (33)
to match the one used for K-means in this library (note that MBKM relies on a different internal stopping condition, despite using the same tolerance parameter; see the scikit-learn documentation).
First, we compared the algorithms’ performance for two real-world datasets included in the scikit-learn library, namely, covtype ($s=54$, p=581 012p=581\,012, $\ell=7$) and kddcup99 ($s=41$, p=494 021p=494\,021, $\ell=5$). For each dataset, we generated 10 sets of initial cluster centers using scikit-learn’s kmeans_plusplus function (the standard warm-start technique used for K-means and MBKM) with different random seeds. Subsequently, we executed K-means, MBKM and SNSM (for $m=0$ and $m=5$) using each of these initializations. The batch size for MBKM was set to 5000.
We used the following three metrics to evaluate clustering quality:
  • The Rand Index (RI), which measures the similarity between the clustering assignments and the ground truth sample labels. A perfect labeling achieves an RI of $1$, whereas poorly agreeing assignments yield lower scores.
  • The Silhouette Coefficient (SC), which for a single data point is given by the ratio $(sc_{1}-sc_{2})/\max\{sc_{1},sc_{2}\}$, where $sc_{1}$ is the mean distance between the data point and all other points in the next nearest cluster, and $sc_{2}$ is the mean distance between the data point and all other points in the same cluster. The overall SC for a dataset is computed as the average across all samples. Higher values indicate better-defined clusters. Due to its computational complexity, the SC is estimated using a random subsample of 50 00050\,000 points.
  • The Calinski–Harabasz index (CH), which is defined as the ratio of the sum of between-clusters dispersion and of within-cluster dispersion. Higher scores indicate dense and well-separated clusters. For further details, we refer the reader to the scikit-learn documentation.
Unlike the RI, neither the SC nor the CH measures require knowledge of the ground truth labels.
Table 4 summarizes the performance of the evaluated solvers. We observe that SNSM exhibits a competitive computational cost compared to the scikit-learn implementations, even outperforming them in some instances. In particular, for the kddcup99 dataset, SNSM attained the lowest running time, while maintaining a clustering quality comparable to that of K-means and significantly better than MBKM in terms of function value and clustering metrics. Regarding clustering quality, SNSM (both $m=0$ and $m=5$) consistently matched or improved upon the results obtained by K-means across the considered metrics, and clearly outperformed MBKM. Recall that MBKM’s random mini-batch sampling mechanism struggles on datasets with high concentration of exact duplicates, such as kddcup99, which may explain the poor values of the objective function and the SC and CH metrics.
Data sets Method Time FunctionvalueFunctionvalue Clustering quality
RI SC CH
[…]
Table 4: Performance comparison between SNSM, K-means, and MBKM for clustering on two real-world datasets
Lastly, we conducted a stress test to evaluate the scalability of SNSM with respect to problem size. We used the make_blobs function from scikit-learn to generate synthetic datasets under three distinct configurations. In each configuration, we varied only a single problem parameter while keeping the others fixed:
  • Data points: For $s=10$ and $\ell=20$, we set $p\in\bigl\{10^{5}\cdot j:1\leq j\leq 20\bigr\}$.
  • Data dimension: For $p=10^{6}$ and $\ell=20$, we set $s\in\bigl\{10^{2}\cdot j:1\leq j\leq 20\bigr\}$.
  • Clusters: For $s=100$ and $p=10^{6}$, we set $\ell\in\bigl\{20\cdot j:1\leq j\leq 20\bigr\}$.
For each configuration, we ran K-means, MBKM, and SNSM ($m=5$) across 10 random seeds. Figure 7.1 presents the average running times on a logarithmic scale. These results highlight SNSM’s strong performance compared to the other two algorithms and indicate that SNSM scales at the same rate as the K-means algorithm for all the tested parameters.
Figure 5: Scalability of SNSM, K-means and MBKM under increasing problem sizes